OpenStudy (anonymous):

Maximize A=xy if 2x+y=16

OpenStudy (anonymous):

soooooo y=16-2x so put that within the equation A=x(16-2x) A=16x-2x^2 taking the derivative of a you get A'(x)=16-4x when A'= 0 you get -16=-4x 4=x if x = 4 y=16-2(4)= 8 x=4 y=8

OpenStudy (anonymous):

you can check if you want if x=1 y =14 A=14 if x=2 y= 12 A=24 ...... on and on

OpenStudy (anonymous):

hope this helps

OpenStudy (anonymous):

=]

OpenStudy (anonymous):

but what is the answer or maximum. I was told to maximize it.

OpenStudy (anonymous):

I also got x=4 and y=8 but I still don't know what the answer is

OpenStudy (anonymous):

how do you get A=24 if x=2??? A= 16-4(2)=2???? no it equals 8, check your math

OpenStudy (anonymous):

2(2)+y=16 4+ y = 16 y=12 so if y = 12 wen x =2, you get xy=2*12 = 24

OpenStudy (anonymous):

answer to maximum is x=4 y = 8

OpenStudy (anonymous):

i'm saying if you aren't sure you can check all the possibles since your number are actually really small it'd be like only 8 or so calculations

OpenStudy (anonymous):

so your max will be 32. my math is correct

OpenStudy (anonymous):

I also got 32